OpenStudy (anonymous):

4-(6x^2-2)/(15)=(5x^2)/(3) - (5(x^2-2))/2

OpenStudy (anonymous):

do you know answer of this?

OpenStudy (anonymous):

@L-Lawliet-L

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

If I would, i would not have asked it.

OpenStudy (anonymous):

ok , then try to solve it. take LCM on both sides.

OpenStudy (anonymous):

\[\frac{ 5x ^{2} }{ 3 }-\frac{ 5x^2-10 }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ 10x^2 }{ 6 }-\frac{ 15x^2-30 }{6 }\]

OpenStudy (anonymous):

LCM on RHS will be 6

OpenStudy (anonymous):

\[\frac{ 10x^2-15x^2-30 }{6}\]

OpenStudy (anonymous):

the equation that you write is divide by 6 and also its not -30 , it will be +30

OpenStudy (anonymous):

solve it. after solving LHS part take to RHS.

OpenStudy (anonymous):

\[4-\frac{ 6x^2-2 }{ 15 }=\frac{ -5x^2+30 }{ 6 }\]

OpenStudy (anonymous):

yes, you are right up to this step.

OpenStudy (anonymous):

now take LCM on LHS first

OpenStudy (anonymous):

\[\frac{ 60 }{ 15 }-\frac{ 6x^2-2 }{ 15 }=\frac{ -5x^2+30 }{ 6 }\]

OpenStudy (anonymous):

solve it and then cross multiply it.

OpenStudy (anonymous):

\[\frac{ 60 }{ 15 }-\frac{ 6x^2-2 }{ 15 }-(\frac{ -5+30 }{ 6 })\]

OpenStudy (anonymous):

now you are doing wrong. 15 is denominator of 60-6x^2+2 whole.

OpenStudy (anonymous):

don't take RHS Part to LHS , you first cross multiply it. it will make it easy.

OpenStudy (anonymous):

@L-Lawliet-L

OpenStudy (anonymous):

??????